An electron is at the origin, and an ion with charge + 5 e is at x = 15nm .

Find a point where the electric field is zero.

x=___ nm
Answer #1

(15 + x) 2
Answer #2


let EF at zero at x unit from origin

here distance between the charges is 15 nm

so at this point of zero EF, EF 1 = EF2

so kq1/x^2 = Kq2/(15+x)^2

15+x/x =   sqrt(5) = 2.236

15+x = 2.236 x

1.236x = 15 nm

x = 12.13 nm
Answer #3

E = F/Q = kQ/r